For the last five fays the average temperature was 63F. If the temperature for the first four days were 56'f,60'f,79'f,70'f what was the temperature for the fifth day? Can you show me how to solve that. Pls. Thank you
Found 2 solutions by stanbon, jorel1380:
Answer by stanbon(75887)

For the last five days the average temperature was 63F. If the temperature for the first four days were 56'f,60'f,79'f,70'f what was the temperature for the fifth day? Can you show me how to solve that.
-------
Average for 5 days = 63 degrees
--------------------------
(56 + 60 + 79 + 70 + x)/5 = 63
-----------------
265+x = 315
x = 50 degrees (temp on the day 5)
-----------
Cheers,
Stan

Answer by jorel1380(3719)

5 days with average temp 63 equals 5*63, or 315 total points.
315-56-60-79-70=50 as the temp for the 5th day. ☺☺☺☺